HUBBLE SPACE TELESCOPE DAILY REPORT #2172 PERIOD COVERED: 0000Z (UTC) 07/23/98 - 0000Z (UTC) 07/24/98 Daily Status Report as of 205/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed NIC/1/2/3 7859 (Blue Compact Dwarf Galaxies -- Young or Old?) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the compact young dwarf galaxy NGC-4214. These blue compact dwarf galaxies challenge models of galaxy formation to explain the presence of small, low-metallicity galaxies experiencing vigorous bursts of star formation. The observations were completed as scheduled, and no problems were reported. 1.3 Completed Nine Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.4 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.6 Completed STIS/CCD 7928 (STIS Spectroscopic CCD Flatfielding -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat field calibration images in the STIS spectroscopic mode. The calibrations were completed as planned, and no problems were reported. 1.7 Completed Four Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.8 Completed NIC/2 7835 (A Search for Superplanets Embedded in Beta Pic & Vega-like Circumstellar Disks) The Near Infrared Camera (Camera Two) was used to make observations of the circumstellar disk around star HD102647 in order to search for superplanets. The discovery of a circumstellar disk orbiting Beta Pictoris as well as the discovery of other similar IR-excess Vega-like systems have provided compelling evidence that our solar system, or at least many of the processes which are thought to have created it, may be far from unique. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.10 Completed STIS/CCD/NIC/2/3/WF/PC-2 7778 (Absorption Lines From Very Highly Ionized Atoms in Front of Quasars) The Space Telescope Imaging Spectrograph (CCD), the WF/PC-2 and the Near Infrared Camera (Cameras Two and Three) were used to make observations of quasar HS1700+641 to study foreground absorption lines from highly ionized atoms. Bright quasars that have a large z{em} and no intervening Lyman-limit systems with large column densities present an excellent opportunity to learn more about distant intergalactic gas regions at very high temperatures. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/MA1 7471 (STIS Snapshot Search for UV-bright High Redshift Quasars) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet observations of the bright high redshifted quasar HS1111+791. This quasar shows no signs of strong absorption in its optical spectra, and its He-II Lambda 304 A0 line is accessible with HST. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed NIC/1/2/3 7855 (The Evolution of Powerful Radio Galaxies) The Near Infrared Camera (Camera One, Two and Three) was used to make observations of the powerful radio galaxy 3C456. Our proposed NICMOS observations combined with our multi-wavelength observing program will finally provide us with the data we need to build a picture of how powerful radio galaxies evolve. The observations were completed as planned, and no porblems were reported. 1.13 Completed NIC/1/2/3 7824 (Deep NIC Images of 20 High-Z Damped Ly-Alpha Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the high-Z damped Lyman-Alpha (DLA) galaxy PKS0458-02. The HST imaging together with ground-based spectroscopy will unify absorption-line studies and faint galaxy imaging studies, providing a comprehensive picture of the history of assembly, star formation, gas consumption, and chemical enrichment in the DLA galaxies. The observations were completed as planned, and no problems were reported. 1.14 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 4 Successful : 4 As detailed in HSTAR 6592 there was a one-minute, 25-second loss of lock beginning at 205/010605Z. One observation in each of two proposals (occurring during the period of the next report) were affected: STIS 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) and NICMOS 7829 (Mapping H_2 Emission around T Tauri Stars) were affected. 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Operations Notes: The SSR EDAC error counter was cleared five times using ROP SR-1. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH